# de Rham affineness of the Nygaard filtered prismatization in positive characteristic

### v1

@@ Abstract.
Let $k$ be a perfect ring of characteristic $p>0$, and let $R$ be an animated $k$-algebra. This note aims to show that the Nygaard filtered prismatization $R^{\mathrm{Nyg}}$ of $R$ is naturally isomorphic, as a stack over $k^{\mathrm{Nyg}}$, to the relative spectrum over $k^{\mathrm{Nyg}}$ of the Rees algebra of the Nygaard filtered prismatic cohomology of $R$ relative to $k$. In doing so, we axiomatise the functorial affineness property displayed by the relative Nygaard filtered prismatization, and dub it de Rham affineness after the fundamental example of the functor sending an animated ring to its relative de Rham stack. While we treat this concept as an organising tool for the author’s forthcoming work studying the syntomification of Frobenius liftable schemes [Sah25c], we are able to frame some questions based on a structural theorem of independent interest: a functor to stacks which is de Rham affine often arises via ring stacks through transmutation.
@@ 1. Introduction
The $E_{\infty}$-ring $\mathrm{dR}_{X/k}$ can also be recovered via the theory of the de Rham stack, which produces, for such $X$, a stack $(X/k)^{\mathrm{dR}}$ with the property that, as $E_{\infty}$-algebras over $k$, one has
@@
Here $\mathrm{DAlg}_{A}$ denotes the -$\infty\text{-category}$ of Bhatt-Mathew and Raksit’s derived $A$-algebras [Rak20, § 4] and $\mathrm{DAlg}_{A}^{(p,I)}$ is the localizing subcategory of $\mathrm{DAlg}_{A}$ consisting of $(p,I)$-complete derived $A$-algebras. The latter category satisfies several pleasing properties and is studied reasonably carefully in [Sah25a, § 2].
@@
More precisely, the relative pushforward of the structure sheaf of $(R/A)^{\mathrm{Nyg}}$ along the structure morphism to $\operatorname{Spf}(A)\times\mathbf{A}^{1}/\mathbf{G}_{m}$ is naturally a filtered $(p,I)$-complete derived algebra under the identification [^3] For all such folklore identifications, we refer the reader to [Sah25a, §§ 2-4]
@@
In this note we specialize to the case when the prism $(A,I)=(W(k),(p))$ and study the relative derived stack $R^{\mathrm{Nyg}}\to k^{\mathrm{Nyg}}$ for any animated $k$-algebra $R$ with the aim of proving an analog of Theorem 1.2 in this setting. We refer the reader to the forthcoming Remark 1.19 on the mutatis mutandis generalization of this result to mixed characteristic pursued in [Sah25b] and to § 1.4 for an explanation of the context .
@@ 1.2. Main result
The theorem remains true mutatis mutandis when one works with relative Nygaard filtered prismatization which is the subject of [Sah25b], as the latter hasn’t been studied in detail before (but see [LM25] and [Pen25] for some definitions). Note that the proof can be easily made to work relative to perfect prisms, but we defer that to [Sah25b].
@@
The proof proceeds by building a natural map and then by reduction the case of $k[x]$ for both the functors in Theorem 1.13. Set $A=k[x]$ and $S=k[x^{1/p^{\infty}}]$. Using the criteria abstracted in Lemma 2.18, it will suffice to show that the natural map $\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{S/k}))\to\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{A/k}))$ is an fpqc epimorphism of sheaves of spaces. This can be obtained from a more concrete flatness statement about the morphism of derived algebras $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{A/k}\to\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{S/k}$, which is obtained via the Rees construction and a constructible stratification flatness argument as explained in Appendix A. The main difficulty then is in setting up a theory of derived algebras on non-trivially formal derived stacks which was accomplished in [Sah25a]. The remaining effort in this paper is explaining the construction of the stack $\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{R/k})$ for an animated $k$-algebra $R.$
@@
This note is second in a series of papers dedicated to studying the syntomification (in the sense of [Bha22]) of Frobenius liftable schemes. The other papers being [Sah25a], [Sah25b] and [Sah25c]. We refer the reader to [Sah25a, § 5] for more context, but we recall here the main context of Theorem 1.13. In [Sah25b] we will show that if $(\mathfrak{S},I)$ is a Breuil-Kisin (or more generally any bounded orientable prism) then the relative Nygaard filtered prismatization $(-/\mathfrak{S})^{\mathrm{Nyg}}$ is de Rham affine in an appropriate sense. In particular when $\mathfrak{S}/I=W(k)$ then the functor sending derived formal $W(k)$-scheme $X\mapsto(X/W(k))^{\mathrm{dR},+}$, the Hodge filtered de Rham stack, is de Rham affine relative to $\mathbf{A}^{1}/\mathbf{G}_{m}$ i.e. a mixed characteristic lift of the statement of Corollary 1.14. Using an observation of [BMS19], this along with Theorem 5.3 will allow us to show in [Sah25c] that if $X$ is a derived $\delta$-scheme over $W(k)$ and $X_{p=0}$ is the derived mod $p$-reduction, then there is a pullback diagram
@@ 1.5. Conventions
The accessibility condition implies that $\mathrm{QCoh}(X),$ defined by right Kan extensions from $\mathrm{DAff}_{X}$, is presentable (for example by [Lur04, Proposition 3.2.2]). Similarly we will be interested in the -$\infty\text{-category}$ $\mathrm{DAlg}(X)$ of derived algebras in quasi-coherent sheaves on a stack $X$. These are defined again by right Kan from $\mathrm{DAff}_{X}$ and are monadic over $\mathrm{QCoh}(X)$ and presentable (by the presentability of $\mathrm{QCoh}(X)$) although we never need the former fact. We will be interested in the classification results for $\mathrm{DAlg}(k^{\mathrm{Nyg}}),\mathrm{DAlg}(\mathbf{A}^{1}/\mathbf{G}_{m})$ and $\mathrm{DAlg}(B\mathbf{G}_{m})$ as explained in [Sah25a, § 4].
@@ 2. Generalities on de Rham affineness
In this section we study some generalities on de Rham affineness as in Definition 1.6 which will serve to axiomatize the notion and streamline the proof. Our viewpoint on this property is that it is a useful organizational tool for our work in the forthcoming [Sah25b] and [Sah25c]. However we are able to show some structural properties like Corollary 2.6 which may be interesting in their own right.
@@ 4. The Rees stack $\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{R/k})$
We recall the construction of the stack in Construction 4.5. But before that we recall the Rees algebra construction for derived algebras, discussed in [Sah25a, § 4], in our context.
@@
If $R$ is an animated ring over $k$, then $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}\in\mathrm{Fil}\mathrm{DAlg}_{\mathrm{Fil}_{(p)}W(k)}^{{p\text{-}\mathrm{comp}}}.$ Indeed, the fact that $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}$ receives an $E_{\infty}$-map from $\mathrm{Fil}_{(p)}W(k)$ is formal. The derived algebra structure can be deduced from the proof of [BL22a, Proposition 5.1.1] and will be explained in slightly more detail in [Sah25b] [^11] We sketch the argument. In the proof of [BL22a, Proposition 5.1.1], the functor $\mathrm{Fil}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{-/k})$ is left and right Kan extended from the category $\mathcal{C}$ of ‘large’ quasisyntomic (hence discrete) $k$-algebras. Here a $k$-algebra $R$ is ‘large’ if it is generated by the image of $R^{\flat}:=\lim_{x\to x^{p}}R$ under the canonical first coordinate map $R^{\flat}\to R$. The values of $\mathrm{Fil}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{-/k})$ on $\mathcal{C}$ is discrete and $\mathrm{Fil}\mathrm{DAlg}_{\mathrm{Fil}_{(p)}W(k)}^{{p\text{-}\mathrm{comp}}}$ is presentable, the left and right Kan extensions are computed in $\mathrm{Fil}\mathrm{DAlg}_{\mathrm{Fil}_{(p)}W(k)}^{{p\text{-}\mathrm{comp}}}$. .
Now as explained in [Sah25a, § 4], there is a chain of equivalences of -$\infty\text{-categories}$
@@
where the right most category denotes the -$\infty\text{-category}$ of $p$-complete graded derived algebras over the graded $p$-complete ring $W(k)\langle u,t\rangle/(ut-p).$ In what follows we identify the left most and right most categories implicitly and denote the image of $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}$ in either as $\mathrm{Rees}(\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}).$ [^12] In [Sah25a] we use the notation $\pi_{!}\mathrm{Rees}(\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k})$ to denote the comodule structure over the coaction given by bicommutative bialgebra $W(k)[\mathbf{Z}]$, in other words, to remember the grading. We drop that notation for now since no confusion can arise.
@@
Fix an animated $k$-algebra $R$. For the various identifications of $\mathrm{DAlg}$ on prestacks we refer the reader to [Sah25a, § 4]. Using the notation in the statement of Lemma 4.2 we have
@@
After Lemma Construction 4.10, Lemma 4.11 and Remark 4.12, this is [Bha22, Theorem 5.5.10 ]. [^14] As written, it seems there is a slight gap in the justification of item (4) of loc.cit. A different proof was communicated to the author by Bhargav Bhatt and will appear in [Sah25b].

### v2

# arXiv:2512.19348v2[math.AG] 28 Feb 2026

@@ Abstract.
Let $k$ be a perfect ring of characteristic $p>0$, and let $R$ be an animated $k$-algebra. This note aims to show that the Nygaard filtered prismatization $R^{\mathrm{Nyg}}$ of $R$ is naturally isomorphic, as a stack over $k^{\mathrm{Nyg}}$, to the relative spectrum over $k^{\mathrm{Nyg}}$ of the Rees algebra of the Nygaard filtered prismatic cohomology of $R$ relative to $k$. In doing so, we axiomatise the functorial affineness property displayed by the relative Nygaard filtered prismatization, and dub it de Rham affineness after the fundamental example of the functor sending an animated ring to its relative de Rham stack. While we treat this concept as an organising tool for the author’s forthcoming work studying the syntomification of Frobenius liftable schemes [Sah25b], we are able to frame some questions based on a structural result of independent interest: a functor to stacks which is de Rham affine often arises via ring stacks through transmutation.
@@ 1. Introduction
The $E_{\infty}$-ring $\mathrm{dR}_{X/k}$ can also be recovered via the theory of the de Rham stack. This produces, for smooth $X$, a stack $(X/k)^{\mathrm{dR}}$ so that as $E_{\infty}$-algebras over $k$, one has
@@
Here $\mathrm{DAlg}_{A}$ denotes the -$\infty\text{-category}$ of Bhatt-Mathew and Raksit’s derived $A$-algebras [Rak20, § 4] and $\mathrm{DAlg}_{A}^{(p,I)}$ is the localizing subcategory of $\mathrm{DAlg}_{A}$ consisting of $(p,I)$-complete derived $A$-algebras. The latter category satisfies several pleasing properties and is studied reasonably carefully in [Sah26, § 2].
@@
More precisely, the relative pushforward of the structure sheaf of $(R/A)^{\mathrm{Nyg}}$ along the structure morphism to $\operatorname{Spf}(A)\times\mathbf{A}^{1}/\mathbf{G}_{m}$ is naturally a filtered $(p,I)$-complete derived algebra under the identification [^3] For all such folklore identifications, we refer the reader to [Sah26, §§ 2-4]
@@
In this note we specialize to the case when the prism $(A,I)=(W(k),(p))$ and study the relative derived stack $R^{\mathrm{Nyg}}\to k^{\mathrm{Nyg}}$ for any animated $k$-algebra $R$ with the aim of proving an analog of Theorem 1.2 in this setting. We refer the reader to the forthcoming Remark 1.19 on the mutatis mutandis generalization of this result to mixed characteristic pursued in [Sah25a] and to § 1.4 for an explanation of the context .
@@ 1.2. Main result
The theorem remains true mutatis mutandis when one works with relative Nygaard filtered prismatization which is the subject of [Sah25a], as the latter hasn’t been studied in detail before (but see [LM25] and [Pen25] for some definitions). Note that the proof can be easily made to work relative to perfect prisms, but we defer that to [Sah25a].
@@
The proof proceeds by building a natural map and then by reduction the case of $k[x]$ for both the functors in Theorem 1.13. Set $A=k[x]$ and $S=k[x^{1/p^{\infty}}]$. Using the criteria abstracted in Lemma 2.18, it will suffice to show that the natural map $\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{S/k}))\to\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{A/k}))$ is an fpqc epimorphism of sheaves of spaces. This can be obtained from a more concrete flatness statement about the morphism of derived algebras $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{A/k}\to\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{S/k}$, which is obtained via the Rees construction and a constructible stratification flatness argument as explained in Appendix A. The main difficulty then is in setting up a theory of derived algebras on non-trivially formal derived stacks which was accomplished in [Sah26]. The remaining effort in this paper is explaining the construction of the stack $\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{R/k})$ for an animated $k$-algebra $R.$
@@
This note is second in a series of papers dedicated to studying the syntomification (in the sense of [Bha22]) of Frobenius liftable schemes. The other papers being [Sah26], [Sah25a] and [Sah25b]. We refer the reader to [Sah26, § 5] for more context, but we recall here the main context of Theorem 1.13. In [Sah25a] we will show that if $(\mathfrak{S},I)$ is a Breuil-Kisin (or more generally any bounded orientable prism) then the relative Nygaard filtered prismatization $(-/\mathfrak{S})^{\mathrm{Nyg}}$ is de Rham affine in an appropriate sense. In particular when $\mathfrak{S}/I=W(k)$ then the functor sending derived formal $W(k)$-scheme $X\mapsto(X/W(k))^{\mathrm{dR},+}$, the Hodge filtered de Rham stack, is de Rham affine relative to $\mathbf{A}^{1}/\mathbf{G}_{m}$ i.e. a mixed characteristic lift of the statement of Corollary 1.14. Using an observation of [BMS19], this along with Theorem 5.3 will allow us to show in [Sah25b] that if $X$ is a derived $\delta$-scheme over $W(k)$ and $X_{p=0}$ is the derived mod $p$-reduction, then there is a pullback diagram
@@ 1.5. Conventions
The accessibility condition implies that $\mathrm{QCoh}(X),$ defined by right Kan extensions from $\mathrm{DAff}_{X}$, is presentable (for example by [Lur04, Proposition 3.2.2]). Similarly we will be interested in the -$\infty\text{-category}$ $\mathrm{DAlg}(X)$ of derived algebras in quasi-coherent sheaves on a stack $X$. These are defined again by right Kan from $\mathrm{DAff}_{X}$ and are monadic over $\mathrm{QCoh}(X)$ and presentable (by the presentability of $\mathrm{QCoh}(X)$) although we never need the former fact. We will be interested in the classification results for $\mathrm{DAlg}(k^{\mathrm{Nyg}}),\mathrm{DAlg}(\mathbf{A}^{1}/\mathbf{G}_{m})$ and $\mathrm{DAlg}(B\mathbf{G}_{m})$ as explained in [Sah26, § 4].
@@ 2. Generalities on de Rham affineness
In this section we study some generalities on de Rham affineness as in Definition 1.6 which will serve to axiomatize the notion and streamline the proof. Our viewpoint on this property is that it is a useful organizational tool for our work in the forthcoming [Sah25a] and [Sah25b]. However we are able to show some structural properties like Corollary 2.6 which may be interesting in their own right.
@@ 4. The Rees stack $\mathcal{R}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{R/k})$
We recall the construction of the stack in Construction 4.5. But before that we recall the Rees algebra construction for derived algebras, discussed in [Sah26, § 4], in our context.
@@
If $R$ is an animated ring over $k$, then $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}\in\mathrm{Fil}\mathrm{DAlg}_{\mathrm{Fil}_{(p)}W(k)}^{{p\text{-}\mathrm{comp}}}.$ Indeed, the fact that $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}$ receives an $E_{\infty}$-map from $\mathrm{Fil}_{(p)}W(k)$ is formal. The derived algebra structure can be deduced from the proof of [BL22a, Proposition 5.1.1] and will be explained in slightly more detail in [Sah25a] [^11] We sketch the argument. In the proof of [BL22a, Proposition 5.1.1], the functor $\mathrm{Fil}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{-/k})$ is left and right Kan extended from the category $\mathcal{C}$ of ‘large’ quasisyntomic (hence discrete) $k$-algebras. Here a $k$-algebra $R$ is ‘large’ if it is generated by the image of $R^{\flat}:=\lim_{x\to x^{p}}R$ under the canonical first coordinate map $R^{\flat}\to R$. The values of $\mathrm{Fil}_{\mathrm{Nyg}}(F^{*}{{\mathbbl{\Delta}}}_{-/k})$ on $\mathcal{C}$ is discrete and $\mathrm{Fil}\mathrm{DAlg}_{\mathrm{Fil}_{(p)}W(k)}^{{p\text{-}\mathrm{comp}}}$ is presentable, the left and right Kan extensions are computed in $\mathrm{Fil}\mathrm{DAlg}_{\mathrm{Fil}_{(p)}W(k)}^{{p\text{-}\mathrm{comp}}}$. .
Now as explained in [Sah26, § 4], there is a chain of equivalences of -$\infty\text{-categories}$
@@
where the right most category denotes the -$\infty\text{-category}$ of $p$-complete graded derived algebras over the graded $p$-complete ring $W(k)\langle u,t\rangle/(ut-p).$ In what follows we identify the left most and right most categories implicitly and denote the image of $\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}$ in either as $\mathrm{Rees}(\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k}).$ [^12] In [Sah26] we use the notation $\pi_{!}\mathrm{Rees}(\mathrm{Fil}_{\mathrm{Nyg}}F^{*}{{\mathbbl{\Delta}}}_{R/k})$ to denote the comodule structure over the coaction given by bicommutative bialgebra $W(k)[\mathbf{Z}]$, in other words, to remember the grading. We drop that notation for now since no confusion can arise.
@@
Fix an animated $k$-algebra $R$. For the various identifications of $\mathrm{DAlg}$ on prestacks we refer the reader to [Sah26, § 4]. Using the notation in the statement of Lemma 4.2 we have
@@
After Lemma Construction 4.10, Lemma 4.11 and Remark 4.12, this is [Bha22, Theorem 5.5.10 ]. [^14] As written, it seems there is a slight gap in the justification of item (4) of loc.cit. A different proof was communicated to the author by Bhargav Bhatt and will appear in [Sah25a].